FEL Upgrade Project Weekly Brief
November 18-27, 2002
This report sums up the last week and half of installation, check-out and commissioning activities on the FEL Upgrade. We had a substantial interruption for most of the project staff last week for the project Semiannual Review on Nov. 20-21. Despite this multi-tasking substantial progress was made on the machine. Highlights include: (1) final assembly of the complete upgrade electron gun. It is now under vacuum with a new cathode ready for the cathode processing and vacuum bakeout procedure; (2) the injector RF is ready for the SRF commissioning; (3) all four optical chicane magnets (DWs) which successfully passed magnetic qualification 2 weeks ago are all installed and aligned in the vault; (4) the first reverse bend dipole (GQ) passed magnetic qualification.

We wish all our collaborators and stakeholders a good Thanksgiving holiday with their friends and family.

The Semiannual Review on Nov. 20-21 was an exhausting event for both the presenters and the reviewers. We appreciate the hard work delivered by all for this important event for the FEL Upgrade project. We look forward to the committee’s report on their technical evaluation of our progress and we appreciate the support shown by both the reviewers and our DOE and DOD program managers for helping us to navigate the FY03 funding issues.

We thank Gil Graf from ONR for expeditiously arranging for the next increment of FY03 funding ($500k) which arrived today, and we thank John Eric from AFRL for a fast start on the process needed for release of the FY03 UV FEL funds.

WBS 4 (Injector):
The two shielded bellows assemblies were pulled, solvent cleaned and vacuumed baked to remove oil contamination (at the 10E-10torr level) that was found during the initial RDGA analysis. The bellows were reinstalled, pumped down and the oil signature disappeared into the noise on the RGA. A wafer was mounted and hydrogen cleaned. The stalk with the wafer was mounted into the gun. On pump down it was discovered the right angle valve had developed a leak across the seat at normal torque values and replaced. The gun is now fully assembled and under vacuum. The external NF3 reservoir and line for wafer conditioning is being worked on. The SF6 tank was fit up to the gun chamber and the operations probe position marked so final machining on the resistor housing could be done.

Gun HVPS - The internal boss to hold the electrical connector for the Conditioning/Running Resistors mechanism was welded by the Machine Shop this week. The tube will be smoothed next week to reduce corona. Final assembly, wiring, and final testing will be done next week.

WBS 6 (RF):
Zone 4 - The download files for this zone were checked. Only a few of the values needed to be corrected. The changes were typically at the 10% level. J. Preble will need to check the cavity Q's to complete the calibration to a reasonable level of accuracy.

Quarter HVPS - Wiring is complete for both power supplies and klystron carts. The last of the LCW plumbing was completed this week. Testing of unit 4's crowbar is being setup for testing today and Monday. Miriam curves will be run Monday. RF into shorts is scheduled for Tuesday. Both units should ready SRF testing Tuesday or Wednesday of next week.

WBS 8 (Instrumentation):
Magnet shunt cable for the 1st, 2nd arc strings, injection dipole, extraction dipole and optical chicane have been pulled into their respective zones. Termination to the shunt chassis is in progress. Work has begun on installing the 200 amp DC power leads for the 1st and 2nd arc magnet strings.

BLM signal and HV cables have been terminated in the MPS rack. Reconfiguration of the MPS rack is in progress. Fabrication of the HV Interlock chassis is complete and functional testing has begun.

Re-termination of the 4 channel BPM drops (1F01 thru 3F11) have been completed. The 25 pin MPS breakout modules were populated. The beamviewer MPS interlock cables were routed between zones and terminated. Cable termination's for the high reflector optical cavity has begun in the vault. Work has begun on installing the magnet shunt cable for the 1st, 2nd Arc strings and optical chicane. The magnet thermal interlock chassis has been installed in Zone 5 and awaiting communications hook up for check out.

Schematic capture for the HV Interlock analog I/O board has been received from Physics and is under revision. EECAD has begun front panel fabrication drawings for the 15 and 25 pin MPS breakout modules. The 1st revision of the BLM system wiring diagram was submitted to EECAD for mark ups. System wiring mark ups for the Magnet Com. upgrade and Magnet strings have been received and under review.

Drive Laser Controls - It was observed that the picomotor that control the drive laser's cavity length was not operating. After investigating the matter, it was found that the multi-axis driver which controls all the picomotors has failed on channel "C". The unit will be sent back to the manufacturer to be serviced. In the mean time, the drive laser's cavity length is being controlled by another channel ("A" or "B") until the unit is pulled-out for the repairs. The final I&C issue on the drive laser controls is to solve a possible failure somewhere in the RS-232 communications with the drive laser control unit. The system operates properly for long periods of time and then behaves unpredictably when the water pump is cycled. We plan to log all serial commands to and from the unit and the ioc to determine whether the problem is software or hardware. If the RS-232 command logger shows nothing unexpected, then the Antares control unit must be considered for service.

Drive Laser Pulse Controller (DLPC) Upgrade - The VME version of the CO305 has checked out well on the bench and is ready to be implemented into iocfel10 in the clean room.

The testing of the optical cavity position sensors (LVDTs) has begun. A bad channel was found on one of the VME cards, it is being sent out for repair (we have a spare). Cabling is complete for the High Reflector end, the motors and encoder (only on 'lift') will be tested as they become available.

Below is the silicon paddle viewer for the strait ahead line spectrometer and for the diverter beamline past the wiggler. The very inexpensive design is compliments of Casey from the machine shop. The assembley simply bolts into one side of a cross and a standard window and camera assembly bolts to the other side. The marks are at 5mm increments, with a viewing surface of 25mm x 100mm.

  WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in
the First Arc.
  • The chamber from the DY is now modified and is with the vacuum group for cleaning and leak testing.
  • DY stand layout is complete that allows the stands to be placed without disturbing the stands already installed for the GY Data is in the hands of the Alignment Group.
  • Power supply for the GY is installed Diversion of the beam after the Wiggler.
  • The layout is complete with hardware and commercial parts well defined.
  • The original IR Demo, DW chamber we are using is in the shop for modification.
  • The stand position for the GW dipole in the hands of the Alignment group

  • Dipoles
    Optical Chicane Dipoles (GW)
  • At MEI, the 6 GW Dipoles for the UV machine (of which we will use one for the Diversion application above) will ship from Oakland on Monday, December 2.

  • Injector Dipoles (DU/DV)
  • No new developments

  • Arc 180 Degree Dipoles (GY)
  • Wang NMR potted the first coil. It will be ready for cure and demolding next week.
  • Wang NMR was successful at gluing the Purcell gap shims on the bottom yoke. They are working on gluing the shims to the upper pole.

  • Arc Bend, Reverse Bend Dipoles (GQ, GX)
  • One GQ has gone through its complete battery of measurements and is near the levels of field accuracy required for qualification. David Douglas has reviewed the data and we are doing a few more tests to refine our understanding.
  • Wang vacuum impregnated the second and third GX coil out of four.

  • Quadrupoles
    Trim Quad (QT)
  • The drawing of the coil (16 required to add the GC vertical corrector function to the trim quad is ready for sign-off.

  • Sextupole (SF)
  • Master Machine completed the remaining "Commissioning Sextupoles" and they were received.

  • Octupole (OT)
  • The specification for the manufacture of these cores is almost complete.

  • Beam Line and Vacuum
  • We have 4 chambers and Master will deliver 3 this week.
  • Chicane chamber fabrication in the Jlab Shop continued with one complete and another being welded with machining on the third done and machining started on the last one.
  • The Linac vacuum is complete to the straight-ahead dump except for the diagnostic viewer added to the branch from the "spectrometer" GW dipole.



    WBS 11 (Optics):
    The week of 11/18 Survey and Alignment aligned the optical cavity vacuum vessels to within ~ 400 microns of the required positions. We can accomplish finer alignment with the adjustments located within the vessel. By Wednesday afternoon today (11/27) the vessels will be connected to the diagnostics crosses holding the ultraviewers and safety shutters.

    The OC side ultraviewer has been supplied with a temporary arrangement that allows it to be actuated. This will aid our alignment efforts. We will follow up with the HR side after the holiday.

    Due to a ~ 2 week delay in Customs, we anticipate receiving the rest of the UHV linear actuators to remotely position the cavity mirrors on Dec 2. This order also includes the actuators for the collimator and most of the optical transport to the optics control room. We received one of the 800 nm HR optics we'll use to set the cavity length.

    We'll sign drawings today (11/27) for the laser safety shutter and the insertable mirror we'll use for wiggler spontaneous emission. At the sametime we will also sign the drawings for the insertable mirror used to route high power beam to the diagnostic beam dump. We are in the process of planning the installation of laser diagnostics in the FEL vault. We received quotes on an absorptive coating for the high power diagnostic beam dump. This process looks promising, but there is some uncertainty on power-handling capability, so we have requested a sample from the vendor. We'll test it in the Mirror Test Stand.

    The vacuum vessels for the high power optical transport are in Procurement, but no action taken as we are awaiting funds. We continue checking drawings for the internal components of mirror cassettes 1,3, &4. Detailing of drawings for the turning cassette 2 internals began. After cleaning and baking, the THz diagnostic was free of oil contamination. The components were reassembled, realigned, and ready for installation.

    We discovered that our inability to control the drive laser cavity length through software was due to the controller. We've contacted the vendor for assistance, but they are closed the entire week. The drive laser developed a small leak in the vicinity of the IR beam shutter, and we repaired it.